OpenStudy (anonymous):

The length of a rectangular piece of tin is twice its width. A two-inch square is cut from each corner so that a box may be formed by folding up the sides. If the box has a volume of 60 cubic inches, what was the width of the original piece of tin? (a) 7" (b) 6" (c) 10" (d) 14" (e) 15"

OpenStudy (mathteacher1729):

How far did you get with this problem before getting stuck?

OpenStudy (anonymous):

i don't even know where to start

OpenStudy (mathteacher1729):

Let's take apart that first sentence... The length of a rectangular piece of tin is twice its width. The length of a rectangular piece of tin = twice its width. The length of a rectangular piece of tin = twice its W. The length of a rectangular piece of tin = 2*W. The L of a rectangular piece of tin = 2*W. L = 2*W

OpenStudy (anonymous):

ok and after the cuts the width is w-4, the length is 2w -4 and the height is 2. Volume = length*width*height so 60 = (2w-4)(w-4)2 and w=7 or w=-1 but a negative width doesn't make sense so the width had to be 7".